HUBBLE SPACE TELESCOPE DAILY REPORT #3091 PERIOD COVERED: 0000Z (UTC) 04/09/02 - 0000Z (UTC) 04/10/02 Daily Status Report as of 100/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.2 Completed Two Sets of ACS/WFC/HRC 8992 (ACS Early Release Observations) The Advanced Camera for Surveys (WFC and HRC) was used to perform early release observations {ERO}, employing a variety of instruments to demonstrate the new capabilities. The observations (not all in this iteration) will include: 1} WFC observations of peculiar galaxies NGC 4676 {The Mice} and UGC 10214. This will be used to study young stellar clusters formed through the process of galaxy interaction. Any dark companion to UGC 10214 will also be searched for in sensitive ACS/WFC observations. 2} Emission line and continuum imaging of the Cone Nebula and M 17 with WFC to demonstrate the improved sensitivity and angular resolution of ACS. 3} Coronagraphic imaging of HD 141569, a main sequence star with circumstellar disk. 4} Coronagraphic imaging with the HRC of the nearby quasar 3C 273 to investigate the physical properties of the host galaxy and environments. There were no reported problems. 1.3 Completed Six Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.4 Completed Two Sets of STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.5 Completed Ten Sets of ACS/WFC/HRC 8948 (CTE Test) The Advanced Camera for Surveys (WFC and HRC) was used to measure the baseline charge transfer efficiency {CTE} performance of the HRC and WFC detectors, emulating the CTE ground calibration testing. Also ,a similar program will perform periodic measurements in order to track the change in the CTE as damage due to radiation exposure accumulates. All observations were successful. 1.6 Completed STIS/CCD 9126 (Galaxy Mass and the Fate of Luminous, Blue Compact Galaxies at z~0.6) The Space Telescope Imaging Spectrograph (CCD) was used to obtain long-slit spectra for a sample of 6 luminous, blue, compact galaxies {LBCGs} at z ~ 0.6. Despite being very luminous, LBCGs have velocity widths Sigma ~ 60 kms and half-light radii r_e ~ 0.5'' {or R_e ~ 3Kpc}. Small sizes and velocity widths suggest LBCGs are low- mass stellar systems , while their blue colors, strong emission lines and low M/L-ratios indicate they are undergoing a major starburst. I f the star-forming process halts after the current burst, models predict that LBCGs will fade by ~2-4 magnitudes after a few Gyrs to reach the low luminosities and surface brightnesses characteristic of spheroidal galaxies. The proposal completed with no reported problems. 1.7 Completed ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.8 Completed ACS/HRC/WFC 9292 (The Nature of Galaxies at z > 4) The Advanced Camera for Surveys (HRC and WFC) was used to further look into recent discoveries of a number of galaxies and quasars at redshifts greater than 5 that has identified the z>5-6 epoch as key to understanding the earliest formation phases for galaxies. However, establishing the characteristics and properties of these earliest galaxies is proving to be a particularly difficult. They are faint, with I{AB} magnitudes around 26-27. Thus, substantial investments of time are needed to obtain high S/N images, while ground-based spectroscopy, even with 8-10 m class telescopes, has provided little more than redshifts. Establishing the physical properties of these galaxies will be a challenge for the foreseeable future. However, there is a subset of this high redshift population that is amenable to more detailed study. These are sources that have been strongly lensed by low redshift clusters. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 11 Successful: 11 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: The NICMOS Cryo Cooler cool down continues nominally. As of yesterday morning, the NICMOS Outlet Temperature (MNPNCOLT) was 84.82K (expected end point 76K), the NCC TA Inlet Temperature (MNPTAILT) was 71.12K (expected end point 68K), the CLI Temperature (MNPLINTT) was 74.36K (expected end point 71K), and the NICMOS Inlet Temperature (MNRNCILT) was 66.37K (expected endpoint 63K). Per an operations request, the NICMOS Cooling System CPL flight software parameters from the 8051 memory were dumped at 100/0200Z. At 100/0202Z, per an operations request, the memory monitor words were set in order to look at DT14. An FOC/ACS status buffer message (54 FOC/ACS 935 P=35, T=46441) was received at 100/060154Z. The software had detected a semaphore timeout. At the time, ACS was in operate mode during a WFC/HRC observation. HSTAR 8601 documents the event. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW